# Coherent Local Explanations for Mathematical Optimization

### v1

@@ Abstract
Machine Learning, ICML, Operations Research, Explainability
@@ 1 Introduction
The field of mathematical optimization plays a crucial role in various domains such as transportation, healthcare, communication, and disaster management (Petropoulos et al., 2024). Since 1940s, significant advancements have been made in this field, leading to the development of complex and effective algorithms like the simplex method and the gradient descent algorithm (Nocedal & Wright, 2006). More recently, the integration of artificial intelligence (AI) and machine learning (ML) techniques has further enhanced optimization methods (Bengio et al., 2021; Scavuzzo et al., 2024).
@@
When using mathematical optimization in practical applications, decision makers must come to a consensus on the main components of the optimization model such as decision variables, objective function, and constraints. They then need to employ an exact or heuristic algorithm to solve the resulting problem. For setting up the model the decision maker has to accurately estimate all necessary parameters for the model and algorithm, e.g., future customer demands or warehouse capacities. However, the solution algorithm can be highly sensitive to even small deviations in these parameters and inaccurate parameter estimations can result in sub-optimal decisions being made.
@@
The analysis of the behavior of an optimization model regarding (small) changes in its problem parameters is widely known as sensitivity analysis (Borgonovo & Plischke, 2016) or parametric optimization (Still, 2018). In both areas, many methods were developed to analyze the model behavior locally and globally, e.g., by one-at-a-time methods, differentiation-based methods or variance-based methods (Borgonovo & Plischke, 2016; Iooss & Lemaître, 2015; Razavi et al., 2021). One of the promising directions mentioned in (Razavi et al., 2021) is the use of ML models to develop sensitivity analysis methods. The main idea of such an approach is to fit an explainable ML model which locally approximates the behavior of the component of the optimization problem to be analyzed (like the optimal objective function value); see e.g., (Wagner, 1995). Usually, linear regression models are fitted because the standardized regression coefficient becomes a natural sensitivity measure. This approach is similar to the LIME method, which is widely used to explain trained ML models (Ribeiro et al., 2016).
@@
Although ML-based sensitivity analysis is effective and model-agnostic, it falls short in providing clear explanations to users when analyzing various components of a model at the same time. Decision makers often need to analyze the main components of an optimization model, such as the objective function value and the values of the decision variables, which are closely intertwined due to the problem’s structure. However, fitting separate linear models to predict the outcome of each component disregards this correlation and leads to incoherent explanations. This can result in situations where either (i) the predicted optimal value does not align with the objective value of the predicted solution, or (ii) the predicted solution violates the constraints of the problem. Inconsistent predictions that do not align with the model’s structure do not enhance understanding of the optimization model; instead, they can cause confusion for the decision makers.
To illustrate this, consider the following simple optimization model
@@
Suppose that the coefficient $a_{12}=4.1$ is the sensitive parameter to analyze. The decision maker is seeking to understand the impact that small changes in this parameter will have on the optimal decision values $x_{1}^{*},x_{2}^{*}$. Fitting two separate linear models on a small number of samples for $a_{12}$ leads to the approximations $x_{1}^{*}\approx 0.11a_{12}$ and $x_{2}^{*}\approx 0.59a_{12}$. If we apply the latter predictions to our nominal parameter value of $a_{12}=4.1$ the constraint value is
@@
In this work, we present a new sampling-based approach called Coherent Local Explanations for Mathematical Optimization (CLEMO). This approach extends the concept of local explanations to multiple components of an optimization model that are coherent with the structure of the model. To incorporate a measure of coherence, we design regularizers evaluating the coherence of the explanation models. We argue that CLEMO is method-agnostic, and hence, it can be used to explain arbitrary exact and heuristic algorithms for solving optimization problems. Lastly, we empirically validate CLEMO on a collection of well-known optimization problems including the shortest path problem, the knapsack problem, and the vehicle routing problem. Our evaluation focuses on accuracy, interpretability, coherence, and stability when subjected to resampling.
Related literature.
To place our work in the literature, we first review the works in explainable AI (XAI) and then move to the studies devoted to explainable optimization.
@@
Recently, there has been a significant amount of research focused on improving the explainability of ML models (Adadi & Berrada, 2018; Bodria et al., 2023; Dwivedi et al., 2023; Linardatos et al., 2021; Minh et al., 2022; Das & Rad, 2020). Common XAI methods include feature-based explanation methods, such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017), and example-based explanations, such as counterfactual explanations, see e.g., the survey by Guidotti (2024). LIME was analyzed and extended in several works regarding its stability (Zhang et al., 2019; Zafar & Khan, 2021) or its use of advanced sampling techniques (Zhou et al., 2021; Saito et al., 2021). In (Dieber & Kirrane, 2020), interviews were conducted with individuals that never worked with LIME before. The research shows that LIME increases model interpretability although the user experience could be improved.
@@
Recently, the notion of explainable and interpretable mathematical optimization attained increasing popularity. Example-based explanation methods such as counterfactuals were introduced to explain optimization models. Korikov et al. (2021) and Korikov & Beck (2023) examine counterfactual explanations for integer problems using inverse optimization. Generalizations of the concept have also been investigated theoretically and experimentally for linear optimization problems (Kurtz et al., 2024). Furthermore, counterfactuals for data-driven optimization were studied in (Forel et al., 2023).
@@
A different approach is incorporating interpretability into the optimization process resulting in intrinsic explainable decision making contrary to the post-hoc explanation method. In (Aigner et al., 2024), for example, the authors study optimization models with an explainability metric added to the objective resulting in a optimization model that makes a trade-off between optimality and explainability. Similarly in (Goerigk & Hartisch, 2023), the authors ensure an interpretable model by using decision trees that resemble the optimization process and hence explain the model by providing optimization rules based on the model parameters.
@@
While -to the best of our knowledge- feature-based explanation methods are scarce for mathematical optimization, parametric optimization and sensitivity analysis are strongly related to these methods. In both fields, the effect of the problem parameters on the model’s output is analyzed where the model’s output can be the optimal value, optimal decision values or even the runtime of the algorithm; see (Still, 2018; Borgonovo & Plischke, 2016; Iooss & Lemaître, 2015; Razavi et al., 2021). Roughly three decades ago Wagner (1995) already presented a global sensitivity method in which he approximated the optimal objective value of linear programming problems such as the knapsack problem with a linear regression model. For this, he used normal perturbations of the model parameters as an input, somewhat a global predecessor of LIME.
@@ 2 Preliminaries
2.1 LIME
Local Interpretable Model-agnostic Explanations (LIME) is an XAI method to produce an explanation for black-box ML models $\bar{h}:\mathcal{Z}\to\mathbb{R}$, which map any data point $\bm{z}$ in the data space $\mathcal{Z}$ to a real value. Given a data point $\bm{z}^{0}$, LIME approximates $\bar{h}$ locally around this point with a surrogate model $\bar{g}$ from a set of explainable models $\mathcal{G}$ (e.g., linear models). To this end LIME samples a set of points $\bm{z}^{1},\ldots,\bm{z}^{N}$ in proximity to $\bm{z}^{0}$ and calculates an optimizer of the problem
$\operatorname*{arg\,min}_{\bar{g}\in\mathcal{G}}\ \sum_{i=0}^{N}w^{i}\ell\left(\bar{g}(\bm{z}^{i}),\bar{h}(\bm{z}^{i})\right)+\Omega(\bar{g}),$ (1)
where $\ell$ is a loss function, $\Omega$ is a complexity measure and $w^{i}$ weighs data points according to their proximity to $\bm{z}^{0}$. LIME uses an indicator function as a complexity measure returning 0 when the number of non-zero features used by $\bar{g}$ is at most $K$ and $\infty$ otherwise. For the weights, LIME uses $w^{i}=\exp(-d(\bm{z}^{i},\bm{z}^{0})^{2}/\nu^{2})$ with distance function $d$ and hyperparameter $\nu$.
2.2 Mathematical Optimization
In mathematical optimization, the aim is to optimize an objective function over a set of feasible solutions. Formally, an optimization problem is given as
@@ 2.2 Mathematical Optimization
where $\bm{x}\in\mathbb{R}^{p}$ are the decision variables, $f$ is an objective function which is parameterized by parameter vector $\bm{\theta}\in\Theta$ and $\mathbb{X}(\bm{\theta})\subseteq\mathbb{R}^{p}$ is the feasible region, again parameterized by $\bm{\theta}$. We call $\bm{\theta}$ the optimization parameters. As an example, one popular class of problems is linear optimization where the problem is defined as $\min\{\bm{c}^{\intercal}\bm{x}:\bm{A}\bm{x}=\bm{b},\bm{x}\geq\bm{0}\}$. In this case we have $\bm{\theta}=(\bm{c},\bm{A},\bm{b})$, $f(\bm{x};\bm{\theta})=\bm{c}^{\intercal}\bm{x}$, and $\mathbb{X}(\bm{\theta})=\{\bm{x}\geq\bm{0}:\bm{A}\bm{x}\geq\bm{b}\}$. The most popular methods to solve linear optimization problems are the simplex method or the interior point method (Bertsimas & Tsitsiklis, 1997).
Many real-world applications from Operations Research involve integer decision variables. In this case the feasible region is given as $\mathbb{X}(\bm{\theta})=\{\bm{x}\in\mathbb{Z}^{p}:\bm{A}\bm{x}\geq\bm{b}\}$. Such so called linear integer optimization problems are widely used, for example for routing problems, scheduling problems and many others (Petropoulos et al., 2024). The most effective exact solution methods are based on branch & bound type algorithms (Wolsey, 2020). However, due to the NP-hardness of this class of problems often large-sized integer problems cannot be solved to optimality in reasonable time. Hence, often problem-specific or general purpose heuristic algorithm are used to quickly calculate possibly non-optimal feasible solutions.
@@ 3 Methodology
In this section, we present CLEMO, a novel method to provide coherent local explanations for multiple components of mathematical optimization problems (2).
Consider a given instance of Problem (2) which is parametrized by $\bm{\theta}^{0}$ which we call the present problem. Additionally, we have a solution algorithm $h$ which we want to explain. The algorithm calculates feasible solutions for every problem instance of Problem (2). Note that this algorithm does not necessarily have to return an optimal solution; our method also works for heuristic or approximation algorithms. The two components we aim to explain in this work are (i) the optimal objective value, and (ii) the values of the decision variables. To this end we fit $p+1$ explainable models combined in the vector-valued function $g:\Theta\to\mathbb{R}^{p+1}$ where $g(\bm{\theta})=(g_{f}(\bm{\theta}),g_{x_{1}}(\bm{\theta}),\dots,g_{x_{p}}(\bm{\theta}))$. Here, $\Theta$ is the parameter space containing all possible parameter vectors $\bm{\theta}$ for (2). For example, the model $g_{x_{i}}$ ideally maps every parameter vector $\bm{\theta}$ to the corresponding solution value of the $i$-th decision variable $x_{i}$ returned by the solution algorithm $h$. For notational convenience, we denote $g(\bm{\theta})=(g_{f}(\bm{\theta}),g_{\bm{x}}(\bm{\theta}))$.
@@
That is, the predictions are aligned with the underlying problem structure. Condition (3) ensures that the predictions for the decision variables $\bm{x}$, when applied to $f$, are leading to the same objective value as the corresponding prediction for the objective value itself. Condition (4) ensures that the predictions of the decision variables are feasible for the corresponding problem.
To find an explanation, we first generate a training data set $\mathcal{D}$ by sampling vectors $\bm{\theta}^{i}\in\Theta$, $i\in[N]_{1}$ which are close to $\bm{\theta}^{0}$. For each problem we apply algorithm $h$ which returns a feasible solution $\bm{x}^{i}$ and the corresponding objective function value $f(\bm{x}^{i},\bm{\theta}^{i})$ for $i\in[N]_{0}$. We denote the returned components of the optimization model by $h(\bm{\theta}^{i}):=(f(\bm{x}^{i};\bm{\theta}^{i}),\bm{x}^{i})$ for $i\in[N]_{0}$. We want to find locally accurate models such that $g_{f}(\bm{\theta}^{i})\approx f(\bm{x}^{i};\bm{\theta}^{i})$ and $g_{\bm{x}}(\bm{\theta}^{i})\approx\bm{x}^{i}$ for all $i\in[N]_{0}$.
Generating Explanations with LIME.
In principle, LIME as in (1) can be applied to any black box function, hence it can be used to explain our solution algorithm $h$. To this end, all explainable predictors in $g$ are fitted by solving the following problem
$\operatorname*{arg\,min}_{g\in\mathcal{G}}~{}\sum_{i=0}^{N}w^{i}\big{(}\ell_{A}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+\Omega(g),$ (5)
where $\mathcal{G}$ contains all $p+1$-dimensional vectors of explainable functions, e.g., linear functions, the scalars $w^{i}\geq 0$ denote the sample weights, $\ell_{A}$ denotes the accuracy loss and $\Omega$ is a complexity measure. If we use linear models $g$, then the corresponding functions $g_{f}$ and $g_{\bm{x}}$ provide explainable predictors for components of the model; i.e., objective value and decision variables. However, this model does not account for the coherence of the calculated predictors with the underlying problem structure (2). As our experiments in Section 4 show, indeed the corresponding predictors in $g$ are usually not coherent, i.e., they violate Conditions 3 and 4 significantly. We use the latter approach as a benchmark method.
Coherent Explanations with CLEMO.
To generate coherent explanations we solve the problem
$\operatorname*{arg\,min}_{g\in\mathcal{G}}~{}\sum_{i=0}^{N}w^{i}\big{(}\ell_{A}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+R_{C}(g(\bm{\theta}^{i}))\big{)},$ (6)
@@
We note that theoretically the coherence conditions could be added as constraints to the minimization problem (5). However, there is no guarantee that a feasible solution $g$ exists, hence we enforce coherence via a regularizer. Similar to LIME, a complexity measure $\Omega$ could be added to the loss function if for example linear models with sparse weights are desired. For ease of notation we omit this term. Note that $\ell_{A}$ and $R_{C}$ can contain hyperparameters to balance all components of the loss function.
@@
$\begin{array}[]{ll}R_{C}(g(\bm{\theta}^{i}))=&\lambda_{C_{1}}(g_{f}(\bm{\theta}^{i})-f(g_{\bm{x}}(\bm{\theta}^{i});\bm{\theta}^{i}))^{2}\\[5.69054pt] &+\lambda_{C_{2}}\delta\left(g_{\bm{x}}(\bm{\theta}^{i}),\mathbb{X}(\bm{\theta}^{i})\right),\end{array}$ (8)
@@
$\delta\left(\bm{x},\mathbb{X}(\bm{\theta}^{i})\right)=\sum_{t=1}^{T}\max\{0,\gamma_{t}(\bm{x},\bm{\theta}^{i})\}.$ (9)
While problem (6) can be applied to different classes of hypothesis sets $\mathcal{G}$, we restrict $\mathcal{G}$ to linear models in this work. In this case we have coefficient vectors $\bm{\beta}_{f},\bm{\beta}_{x_{1}},\dots,\bm{\beta}_{x_{p}}$, and
$g_{f}(\bm{\theta}^{i};\bm{\beta}):=\bm{\beta}_{f}^{\intercal}\bm{\theta}^{i}\text{ and }g_{\bm{x}}(\bm{\theta}^{i};\bm{\beta}):=(\bm{\beta}_{x_{1}}^{\intercal}\bm{\theta}^{i},\dots,\bm{\beta}_{x_{p}}^{\intercal}\bm{\theta}^{i}).$
Problem (6) then becomes
$\operatorname*{arg\,min}_{\bm{\beta}}~{}\sum_{i=0}^{N}w^{i}\big{(}\ell_{A}(\bm{\beta}^{\intercal}\bm{\theta}^{i},h(\bm{\theta}^{i}))+R_{C}(\bm{\beta}^{\intercal}\bm{\theta}^{i})\big{)},$ (10)
where $\bm{\beta}\equiv(\bm{\beta}_{f},\bm{\beta}_{x_{1}},\dots,\bm{\beta}_{x_{p}})^{\intercal}$.
@@
Since we use the squared loss (7) in Problem (10), the first term corresponding to the accuracy loss becomes a convex function of $\bm{\beta}$. For the coherence regularizer the following holds.

### v2

@@ 1 Introduction
The field of mathematical optimization plays a crucial role in various domains such as transportation, healthcare, communication, and disaster management [27]. Since 1940s, significant advancements have been made in this field, leading to the development of complex and effective algorithms like the simplex method and the gradient descent algorithm [26]. More recently, the integration of artificial intelligence (AI) and machine learning (ML) techniques has further enhanced optimization methods [3, 32].
@@
When using mathematical optimization in practical applications, decision makers must come to a consensus on the main components of the optimization model such as decision variables, objective function, and constraints. Afterwards, they need to employ an exact or heuristic algorithm to solve the resulting problem. For setting up the model, the decision maker has to accurately estimate all necessary parameters for the model and algorithm, e.g., future customer demands or warehouse capacities. However, the solution algorithm can be highly sensitive to even small deviations in these parameters and inaccurate parameter estimations can result in sub-optimal decisions being made.
@@
The analysis of the behavior of an optimization model regarding (small) changes in its problem parameters is widely known as sensitivity analysis [6] or parametric optimization [34]. In both areas, many methods were developed to analyze the model behavior locally and globally, e.g., by one-at-a-time methods, differentiation-based methods or variance-based methods [6, 18, 29]. One of the promising directions mentioned in [29] is the use of ML models to develop sensitivity analysis methods. The main idea of such an approach is to fit an explainable ML model which locally approximates the behavior of the component of the optimization problem to be analyzed (like the optimal objective function value); see e.g., [36]. Usually, linear regression models are fitted because the standardized regression coefficient becomes a natural sensitivity measure. This approach is similar to the LIME method, which is widely used to explain trained ML models [30].
@@
Although ML-based sensitivity analysis is effective and model-agnostic, it falls short in providing clear explanations to users when analyzing various components of a model at the same time. Decision makers often need to analyze the main components of an optimization model, such as the objective function value and the values of the decision variables, which are closely intertwined due to the problem’s structure. However, fitting separate linear models to predict the outcome of each component disregards this correlation and leads to incoherent explanations. This can result in situations where either (i) the predicted optimal value does not align with the objective value of the predicted solution, or (ii) the predicted solution violates the constraints of the problem. Inconsistent predictions that do not align with the model’s structure do not enhance understanding of the optimization model; instead, they can cause confusion for the decision makers. To illustrate this, consider the following simple optimization model
@@
Suppose that the coefficient $a_{12}=4.1$ is the sensitive parameter to analyze. The decision maker is seeking to understand the impact that small changes in this parameter will have on the optimal decision values $x_{1}^{*},x_{2}^{*}$. Fitting two separate linear models on a small number of samples for $a_{12}$ leads to the approximations $x_{1}^{*}\approx 0.11a_{12}$ and $x_{2}^{*}\approx 0.59a_{12}$. If we apply the latter predictions to our nominal parameter value of $a_{12}=4.1$, then the constraint value becomes
@@
Contributions. In this work, we present a new sampling-based approach called Coherent Local Explanations for Mathematical Optimization (CLEMO). This approach extends the concept of local explanations to multiple components of an optimization model that are coherent with the structure of the model. To incorporate a measure of coherence, we design regularizers evaluating the coherence of the explanation models. We argue that CLEMO is method-agnostic, and hence, it can be used to explain arbitrary exact and heuristic algorithms for solving optimization problems. Lastly, we empirically validate CLEMO on a collection of well-known optimization problems including the shortest path problem, the knapsack problem, and the vehicle routing problem. Our evaluation focuses on accuracy, interpretability, coherence, and stability when subjected to resampling.
@@
Related literature. Recently, there has been a significant amount of research focused on improving the explainability of ML models [1, 5, 11, 23, 25, 9]. Common XAI methods include feature-based explanation methods, such as LIME [30] and SHAP [24], and example-based explanations, such as counterfactual explanations, see e.g., the survey by Guidotti [16]. LIME was analyzed and extended in several works regarding its stability [39, 38] or its use of advanced sampling techniques [40, 31]. In [10], interviews were conducted with individuals that never worked with LIME before. The research shows that LIME increases model interpretability although the user experience could be improved.
@@
Recently, the notion of explainable and interpretable mathematical optimization attained increasing popularity. Example-based explanation methods such as counterfactuals were introduced to explain optimization models. Korikov et al. [21] and Korikov and Beck [20] examine counterfactual explanations for integer problems using inverse optimization. Generalizations of the concept have also been investigated theoretically and experimentally for linear optimization problems [22]. Furthermore, counterfactuals for data-driven optimization were studied in [12].
@@
A different approach is incorporating interpretability into the optimization process resulting in intrinsic explainable decision making contrary to the post-hoc explanation method. In [2], for example, the authors study optimization models with an explainability metric added to the objective resulting in a optimization model that makes a trade-off between optimality and explainability. Similarly in [15], the authors ensure an interpretable model by using decision trees that resemble the optimization process and hence explain the model by providing optimization rules based on the model parameters.
@@
While -to the best of our knowledge- feature-based explanation methods are scarce for mathematical optimization, parametric optimization and sensitivity analysis are strongly related to these methods. In both fields, the effect of the problem parameters on the model’s output is analyzed where the model’s output can be the optimal value, optimal decision values or even the runtime of the algorithm; see [34, 6, 18, 29]. Three decades ago, Wagner [36] already presented a global sensitivity method in which he approximated the optimal objective value of linear programming problems like the knapsack problem with a linear regression model. For this, he used normal perturbations of the model parameters as an input, somewhat a global predecessor of LIME.
@@ 2 Preliminaries
LIME. Local Interpretable Model-agnostic Explanations (LIME) is an XAI method to produce an explanation for black-box ML models $\bar{h}:\mathcal{Z}\to\mathbb{R}$, which map any data point $\bm{z}$ in the data space $\mathcal{Z}$ to a real value. Given a data point $\bm{z}^{0}$, LIME approximates $\bar{h}$ locally around this point with a surrogate model $\bar{g}$ from a set of explainable models $\mathcal{G}$ (e.g., linear models). To this end, LIME samples a set of points $\bm{z}^{1},\ldots,\bm{z}^{N}$ in proximity to $\bm{z}^{0}$ and calculates an optimizer of the problem
$\operatorname*{arg\,min}_{\bar{g}\in\mathcal{G}}\ \sum\nolimits_{i=0}^{N}w^{i}\ell\left(\bar{g}(\bm{z}^{i}),\bar{h}(\bm{z}^{i})\right)+\Omega(\bar{g}),$ (1)
where $\ell$ is a loss function, $\Omega$ is a complexity measure and $w^{i}$ weighs data points according to their proximity to $\bm{z}^{0}$. LIME uses an indicator function as a complexity measure returning 0 when the number of non-zero features used by $\bar{g}$ is at most $K$, and $\infty$, otherwise. For the weights, LIME uses $w^{i}=\exp(-d(\bm{z}^{i},\bm{z}^{0})^{2}/\nu^{2})$ with distance function $d$ and hyperparameter $\nu$.
Mathematical Optimization. In mathematical optimization, the aim is to optimize an objective function over a set of feasible solutions. Formally, an optimization problem is given as
@@
where $\bm{x}\in\mathbb{R}^{p}$ are the decision variables, $f$ is an objective function which is parameterized by parameter vector $\bm{\theta}\in\Theta$ and $\mathbb{X}(\bm{\theta})\subseteq\mathbb{R}^{p}$ is the feasible region, again parameterized by $\bm{\theta}$. We call $\bm{\theta}$ the optimization parameters. As an example, one popular class of problems belongs linear optimization, where the problem is defined as $\min\{\bm{c}^{\intercal}\bm{x}:\bm{A}\bm{x}=\bm{b},\bm{x}\geq\bm{0}\}$. In this case, we have $\bm{\theta}=(\bm{c},\bm{A},\bm{b})$, $f(\bm{x};\bm{\theta})=\bm{c}^{\intercal}\bm{x}$, and $\mathbb{X}(\bm{\theta})=\{\bm{x}\geq\bm{0}:\bm{A}\bm{x}\geq\bm{b}\}$. The most popular methods to solve linear optimization problems are the simplex method or the interior point method [4].
Many real-world applications from operations research involve integer decision variables. In this case the feasible region is given as $\mathbb{X}(\bm{\theta})=\{\bm{x}\in\mathbb{Z}^{p}:\bm{A}\bm{x}\geq\bm{b}\}$. Such so called linear integer optimization problems are widely used, for example for routing problems, scheduling problems and many others [27]. The most effective exact solution methods are based on branch & bound type algorithms [37]. However, due to the NP-hardness of this class of problems often large-sized integer problems cannot be solved to optimality in reasonable time. Hence, often problem-specific or general purpose heuristic algorithm are used to quickly calculate possibly non-optimal feasible solutions.
@@ 3 Methodology
In this section, we present CLEMO, a novel method to provide coherent local explanations for multiple components of mathematical optimization problems (2). Consider a given instance of Problem, (2) which is parametrized by $\bm{\theta}^{0}$, and we call it the present problem. Additionally, we have a solution algorithm $h$ that we want to explain. The algorithm calculates feasible solutions for every problem instance of Problem (2). Note that this algorithm does not necessarily have to return an optimal solution, since our method also works for heuristic or approximation algorithms. The two components we aim to explain in this work are (i) the optimal objective value, and (ii) the values of the decision variables. To this end, we fit $p+1$ explainable models combined in the vector-valued function $g:\Theta\to\mathbb{R}^{p+1}$ where $g(\bm{\theta})=(g_{f}(\bm{\theta}),g_{x_{1}}(\bm{\theta}),\dots,g_{x_{p}}(\bm{\theta}))$. Here, $\Theta$ is the parameter space containing all possible parameter vectors $\bm{\theta}$ for (2). For example, the model $g_{x_{i}}$ ideally maps every parameter vector $\bm{\theta}$ to the corresponding solution value of the $i$-th decision variable $x_{i}$ returned by the solution algorithm $h$. For notational convenience, we denote $g(\bm{\theta})=(g_{f}(\bm{\theta}),g_{\bm{x}}(\bm{\theta}))$.
@@
That is, the predictions are aligned with the underlying problem structure. Condition (3) ensures that the predictions for the decision variables $\bm{x}$, when applied to $f$, lead to the same objective value as the corresponding prediction for the objective value itself. Condition (4) ensures that the predictions of the decision variables are feasible for the corresponding problem.
To find an explanation, we first generate a training data set $\mathcal{D}$ by sampling vectors $\bm{\theta}^{i}\in\Theta$, $i\in[N]_{1}$ which are close to $\bm{\theta}^{0}$. For each problem, we apply algorithm $h$ which returns a feasible solution $\bm{x}^{i}$ and the corresponding objective function value $f(\bm{x}^{i},\bm{\theta}^{i})$ for $i\in[N]_{0}$. We denote the returned components of the optimization model by $h(\bm{\theta}^{i}):=(f(\bm{x}^{i};\bm{\theta}^{i}),\bm{x}^{i})$ for $i\in[N]_{0}$. We want to find locally accurate models such that $g_{f}(\bm{\theta}^{i})\approx f(\bm{x}^{i};\bm{\theta}^{i})$ and $g_{\bm{x}}(\bm{\theta}^{i})\approx\bm{x}^{i}$ for all $i\in[N]_{0}$.
Generating Explanations with LIME. In principle, LIME as in (1) can be applied to any black box function, hence it can be used to explain our solution algorithm $h$. To this end, all explainable predictors in $g$ are fitted by solving the following problem
$\operatorname*{arg\,min}_{g\in\mathcal{G}}~{}\sum\nolimits_{i=0}^{N}w^{i}\big{(}\ell_{A}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+\Omega(g),$ (5)
where $\mathcal{G}$ contains all $p+1$-dimensional vectors of explainable functions, e.g., linear functions, the scalars $w^{i}\geq 0$ denote the sample weights, $\ell_{A}$ denotes the accuracy loss, and $\Omega$ is a complexity measure. If we use linear models for $g$, then the corresponding functions $g_{f}$ and $g_{\bm{x}}$ provide explainable predictors for components of the model; i.e., objective value and decision variables. However, this model does not account for the coherence of the calculated predictors with the underlying problem structure (2). As our experiments in Section 4 show, indeed the corresponding predictors in $g$ are usually not coherent, i.e., they violate Conditions 3 and 4 significantly. We use the latter approach as a benchmark method.
Coherent Explanations with CLEMO. To generate coherent explanations we solve the problem
$\operatorname*{arg\,min}_{g\in\mathcal{G}}~{}\sum\nolimits_{i=0}^{N}w^{i}\big{(}\ell_{A}(g(\bm{\theta}^{i}),h(\bm{\theta}^{i}))+R_{C}(g(\bm{\theta}^{i}))\big{)},$ (6)
@@
We note that theoretically, the coherence conditions could be added as constraints to the minimization problem (5). However, there is no guarantee that a feasible solution $g$ exists, hence we enforce coherence via a regularizer. Similar to LIME, a complexity measure $\Omega$ could be added to the loss function if, for example, linear models with sparse weights are desired. For ease of notation, we omit this term. Note that $\ell_{A}$ and $R_{C}$ can contain hyperparameters to balance all components of the loss function.
@@
$\begin{array}[]{ll}R_{C}(g(\bm{\theta}^{i}))=&\lambda_{C_{1}}(g_{f}(\bm{\theta}^{i})-f(g_{\bm{x}}(\bm{\theta}^{i});\bm{\theta}^{i}))^{2}+\lambda_{C_{2}}\delta\left(g_{\bm{x}}(\bm{\theta}^{i}),\mathbb{X}(\bm{\theta}^{i})\right),\end{array}$ (8)
@@
$\delta\left(\bm{x},\mathbb{X}(\bm{\theta}^{i})\right)=\sum\nolimits_{t=1}^{T}\max\{0,\gamma_{t}(\bm{x},\bm{\theta}^{i})\}.$ (9)
While problem (6) can be applied to different classes of hypothesis sets $\mathcal{G}$, we restrict $\mathcal{G}$ to linear models in this work. In this case we have coefficient vectors $\bm{\beta}_{f},\bm{\beta}_{x_{1}},\dots,\bm{\beta}_{x_{p}}$, such that $g_{f}(\bm{\theta}^{i};\bm{\beta}):=\bm{\beta}_{f}^{\intercal}\bm{\theta}^{i}$ and $g_{\bm{x}}(\bm{\theta}^{i};\bm{\beta}):=(\bm{\beta}_{x_{1}}^{\intercal}\bm{\theta}^{i},\dots,\bm{\beta}_{x_{p}}^{\intercal}\bm{\theta}^{i})$. For $\bm{\beta}\equiv(\bm{\beta}_{f},\bm{\beta}_{x_{1}},\dots,\bm{\beta}_{x_{p}})^{\intercal}$, Problem (6) then becomes
$\operatorname*{arg\,min}_{\bm{\beta}}~{}\sum\nolimits_{i=0}^{N}w^{i}\big{(}\ell_{A}(\bm{\beta}^{\intercal}\bm{\theta}^{i},h(\bm{\theta}^{i}))+R_{C}(\bm{\beta}^{\intercal}\bm{\theta}^{i})\big{)}.$ (10)
@@
Since we use the squared loss (7) in Problem (10), the first term corresponding to the accuracy loss becomes a convex function of $\bm{\beta}$. For the coherence regularizer, the following holds.